### 12 Feb 2024 v3

# Universality of almost periodicity in bounded discrete time series

Abstract.
We consider bounded discrete time series. From its statistical feature, without any use of the Fourier transform, we find a suitable almost periodic function which approximates the corresponding time series in a local time interval.
Key words and phrases:
almost periodic function, chaos in discrete time, recurrent neural network
2020 Mathematics Subject Classification:
Primary 42A20; Secondary 11B50
1. Introduction
In this paper we consider universality of almost periodicity. More precisely we show that, without any use of the Fourier transform, we find a suitable almost periodic function which approximates the corresponding bounded discrete time series in a local time interval.
Let $y:\mathbb{Z}\to[-1,1]$ be a bounded discrete time series, and first let us discretize the range $[-1,1]$ as follows: For $K\in\mathbb{Z}_{\geq 1}$, we choose $\{a_{k}^{K}\}_{k=0}^{K+1}$ such that (for a technical reason, we employ transcendental numbers)
  • $\{a_{1}^{K},a_{2}^{K},\cdots,a^{K}_{K-1},a_{K}^{K}\}\subset\{\pm e^{-\frac{n}{m}}\ ;m,n\in\mathbb{Z}_{\geq 1}\}\subset\mathbb{R}\setminus\mathbb{Q}$,
  • $-1=a_{0}^{K}<a_{1}^{K}<a_{2}^{K}<\cdots<a_{K+1}^{K}=1$,
  • $\sup_{1\leq k\leq K+1}|a_{k-1}^{K}-a_{k}^{K}|\leq^{\exists}\!\!C/K$.
We discretize $y$ as follows:
$\bar{y}^{K}(t):=\operatorname*{arg\,min}_{a\in\{a^{K}_{k}\}_{k=1}^{K}}|y(t)-(a-0)|,$
where $a-0:=a-\varepsilon$ for any sufficiently small $\varepsilon>0$. In what follows we sometimes omit the index $K$ unless confusion occurs. To define a suitable statistical feature, we need to classify patterns in $\bar{y}$. Let $\sigma_{n}$ ($n=1,2,\cdots,N$) be a permutation operator, namely,
$\sigma_{n}:\{1,2,\cdots,L\}\to\{a_{1},a_{2},\cdots,a_{K-1},a_{K}\}$
($\ell\mapsto\sigma_{n}(\ell)$) with $\sigma_{n}\not=\sigma_{n^{\prime}}$ ($n\not=n^{\prime}$), and we impose the following two conditions for determining $N$:
(1) $\begin{cases}\text{For any $t\in\mathbb{Z}$, there is $n\in\{1,\cdots,N\}$ such that}\\ \qquad\text{$\sigma_{n}(\ell)=\bar{y}(t-\ell)$ for $\ell=1,2,\cdots,L$},\\ \text{For any $n\in\{1,\cdots,N\}$ there is $t\in\mathbb{Z}$ such that}\\ \qquad\text{$\sigma_{n}(\ell)=\bar{y}(t-\ell)$ for $\ell=1,2,\cdots,L$.}\end{cases}$
Note that $N\leq K^{L}$ due to the sequence with repetition. The crucial statistical feature is the following: for any $y:\mathbb{Z}\to[-1,1]$, we have the following conditional probability: there exists $\delta^{K}\in[0,1)$ (depending on $K$) such that for any $n\in\{1,2,\cdots,N\}$,
(2) $P\bigg{(}a_{k(n)}=\bar{y}(\cdot)\bigg{|}\sigma_{n}(\ell)=y(\cdot-\ell)\quad\text{for}\quad\ell=1,2,\cdots,L\bigg{)}\geq 1-\delta^{K}$
with appropriately chosen $a_{k(n)}$ satisfying the following property:
  • Magnitude relation: For any $K^{\prime}>0$ and $t\in\mathbb{Z}$,
    (3) $\left|a^{K}_{k(n_{t}^{K})}-a^{K+K^{\prime}}_{k(n_{t}^{K+K^{\prime}})}\right|<\frac{C}{K},$
    where $n_{t}^{K}$ is determined as
    (4) $\bar{y}^{K}(t-\ell)=\sigma_{n_{t}^{K}}(\ell)\quad\text{for}\quad\ell=1,2,\cdots,L.$
Then we can state the main theorem as follows:
Theorem 1.
Assume $(\sigma_{n},a_{k(n)})$ satisfies the following:
  • (Uniqueness in backward direction.) For any $n^{\prime}\in\{1,2,\cdots,N\}$, there exists at most one $n\in\{1,2,\cdots,N\}$ such that
    (5) $(a_{k(n)},\sigma_{n}(1),\cdots,\sigma_{n}(L-1))=(\sigma_{n^{\prime}}(1),\sigma_{n^{\prime}}(2),\cdots,\sigma_{n^{\prime}}(L)).$
Then, for any $t_{0}\in\mathbb{Z}$, there exists an almost periodic function $u$ such that
(6) $\prod_{t^{\prime}=1}^{t}P\bigg{(}|u(t_{0}+t^{\prime})-y(t_{0}+t^{\prime})|<\frac{C}{K}\bigg{)}\geq\left(1-\delta^{K}\right)^{t}$
for any $t\in\mathbb{Z}_{\geq 1}$.
Remark 1.
Deriving the inequality (6) itself is rather obvious and quite natural due to (2). The nontrivial discovery is that $u$ is a discretized almost periodic function provided by (5), without any use of Fourier transform.
2. Proof of main theorem
Let $\sigma^{*}_{n}$ ($n=1,2,\cdots,N$) be an adjoint type of permutation operator, more precisely, let
(7) $\sigma^{*}_{n}(\ell-1):=\frac{1}{\sigma_{n}(\ell)}$
for $\ell\in\{1,\cdots,L\}$ and $n\in\{1,\cdots N\}$, where $\sigma^{*}_{i}(0):=\sigma_{i}^{*}(L)$. The definition of the following $N\times N$ matrix $X$ is the key to prove the main theorem:
$X:=\left[h\left(\sum_{\ell=1}^{L}\sigma^{*}_{i}(\ell-1)\sigma_{j}(\ell)\right)\right]_{i,j}$
for some $h:\mathbb{R}\to[-1,1]$,
Lemma 2.
$X$ is a regular matrix provided by appropriately chosen $h:\mathbb{R}\to[-1,1]$.
Proof.
First, by the Lindemann–Weierstrass theorem, we see that
(8) $\frac{a_{k^{\prime}}}{a_{k}}\in(\mathbb{R}\setminus\mathbb{Q})\cup\{-1\}\quad(k\not=k^{\prime}),$
(9) $\sum_{\ell=1}^{L}\frac{a_{k^{\prime}_{\ell}}}{a_{k_{\ell}}}\in(\mathbb{R}\setminus\mathbb{Q})\cup\{-L,-L+1,\cdots,-1,0,1,\cdots,L-1\}$
except for the $k_{1}=k^{\prime}_{1},\ k_{2}=k^{\prime}_{2},\cdots,k_{L}=k^{\prime}_{L}$ case. Also
$\sum_{\ell=1}^{L}\frac{a_{k^{\prime}_{\ell}}^{K}}{a_{k_{\ell}}^{K}}=L$
if and only if $k_{1}=k^{\prime}_{1},\ k_{2}=k^{\prime}_{2},\cdots,k_{L}=k^{\prime}_{L}$. By (7), (8), (9) and Lindemann–Weierstrass theorem, we see
$\sum_{\ell=1}^{L}\sigma^{*}_{j}(\ell-1)\sigma_{j}(\ell)=L\quad\text{and}\quad\sum_{\ell=1}^{L}\sigma^{*}_{i}(\ell-1)\sigma_{j}(\ell)\not=L\quad(i\not=j).$
In order to construct an appropriate $h$, we use a finite set $G$ as follows:
$G:=\left\{\sum_{\ell=1}^{L}\sigma^{*}_{i}(\ell-1)\sigma_{j}(\ell):i,j\in\{1,2,\cdots,N\}\right\}\subset\mathbb{R}.$
Note that $L\in G$. Now we take a smooth function $h:\mathbb{R}\to[-1,1]$ satisfying the following:
$\begin{split}&h(L)\in\mathbb{Q}\setminus\{0\},\\ &h(\gamma)\in\left(\{\pm e^{-\frac{n}{m}}\ ;m,n\in\mathbb{Z}_{\geq 1}\}\cup\{0\}\right)\subset(\mathbb{R}\setminus\mathbb{Q})\cup\{0\},\quad\text{for}\quad\gamma\subset G\setminus\{1\}.\end{split}$
Then we can easily check that (applying the Lindemann–Weierstrass theorem)
  • $h(\gamma_{1})h(\gamma_{2})\cdots h(\gamma_{N})\in(\mathbb{R}\setminus\mathbb{Q})\cup\{0\}$  for  $\{\gamma_{n}\}_{n=1}^{N}\in G^{N}\setminus\{\underbrace{1,1,\cdots,1}_{N}\}$,
  • for any $\{\tau_{n^{\prime}}\}_{n^{\prime}=1}^{N!-1}\in\{-1,1\}^{N!-1}$ and $\{\{\gamma_{n,n^{\prime}}\}_{n=1}^{N}\}_{n^{\prime}=1}^{N!-1}\subset G^{N}\setminus\{\underbrace{1,1,\cdots,1}_{N}\}$,
    $\sum_{n^{\prime}=1}^{N!-1}\tau_{n^{\prime}}h(\gamma_{1,n^{\prime}})h(\gamma_{2,n^{\prime}})\cdots h(\gamma_{N,n^{\prime}})\in(\mathbb{R}\setminus\mathbb{Q})\cup\{0\}.$
By applying the above properties, then we see that the determinant of the matrix $X$ is nonzero, since it is expressed as
$|X|=\eta_{1}+\eta_{2},\quad\eta_{1}\in\mathbb{Q}\setminus\{0\},\quad\eta_{2}\in(\mathbb{R}\setminus\mathbb{Q})\cup\{0\}.$
Now we resume the proof of the main theorem. Let us define $N\times N$ matrices: $Y$ and $W$ as follows:
(10) $Y:=\left[\sum_{\ell=1}^{L}\sigma_{i}^{*}(\ell)\sigma_{j}(\ell)\right]_{i,j}\quad\text{and}\quad W:=YX^{-1}.$
By (1), for each $t\in\mathbb{Z}$, there exists a unique column vector
$e_{n_{t}}:=(\underbrace{\overbrace{0,0,\cdots,0,1}^{n_{t}},0\cdots,0}_{N})^{T}$
such that
(11) $\bar{y}(t-\ell)=\sigma(\ell)e_{n_{t}}\quad\text{for}\quad\ell=1,2,\cdots,L,$
where
$\sigma(\ell):=(\sigma_{1}(\ell),\sigma_{2}(\ell),\cdots,\sigma_{N}(\ell)).$
Thus we have
$W\left(h\left(\sum_{\ell=1}^{L}\sigma^{*}(\ell-1)\bar{y}(t-\ell)\right)\right)=WXe_{n_{t}}=Ye_{n_{t}}=\sum_{\ell=1}^{L}\sigma^{*}(\ell)\bar{y}(t-\ell),$
where $\sigma^{*}(\ell):=(\sigma^{*}_{1}(\ell),\sigma^{*}_{2}(\ell),\cdots,\sigma_{N}^{*}(\ell))^{T}$. To find an appropriate almost periodic function $u$, we create the explicit formulas $\{\bar{u}(t)\}_{t\geq t_{0}}$ inductively from initial data $\{\bar{y}(t_{0}-\ell)\}_{\ell=1}^{L}$:
(12) $\begin{split}&\begin{cases}\bar{r}(t_{0})&=h(\sum_{\ell=1}^{L}\sigma^{*}(\ell)\bar{y}(t_{0}-\ell)),\\ \bar{u}(t_{0})&=W^{out}\bar{r}(t_{0}),\end{cases}\\ &\begin{cases}\bar{r}(t)&=h(W\bar{r}(t-1)+\sigma^{*}(L)\bar{u}(t-1)-\sigma^{*}(L)\bar{u}(t-L-1)),\\ \bar{u}(t)&=W^{out}\bar{r}(t),\end{cases}\end{split}$
for $t=t_{0}+1,t_{0}+2,\cdots$, where $W^{out}:=AX^{-1}$ and $A:=(a_{k(1)},a_{k(2)},\cdots,a_{k(N)})$. Rigorously we regard $\bar{u}(t-L-1)$ as $\bar{y}(t-L-1)$ for $t=t_{0}+1,t_{0}+2\cdots,t_{0}+L$. Now we show that $\bar{u}^{K}(t)$ is a discretized periodic function for any $K$. By (2), we have a chain of the permutation operator $(\sigma^{K}_{j},a^{K}_{k(j)})$ as follows:
$(\sigma_{j^{K}_{1}},a^{K}_{k(j_{1})})\rightarrow(\sigma_{j^{K}_{2}},a^{K}_{k(j_{2})})\rightarrow(\sigma_{j^{K}_{3}},a^{K}_{k(j_{3})})\rightarrow\cdots.$
Note that the sequence $j_{1},j_{2},j_{3}\cdots$ is different from $n_{1},n_{2},n_{3}\cdots$ in (3). Since the number of permutation operators $\{\sigma_{n}\}_{n=1}^{N}$ is finite and by backward uniqueness (5), then there is $m\in\mathbb{Z}_{\geq 2}$ such that $(\sigma_{j^{K}_{1}}a_{k(j_{1})}^{K})=(\sigma_{j^{K}_{m}},a_{k(j_{m})}^{K})$. This means that we have a periodic chain for any $K$. Moreover $\lim_{K\to\infty}\bar{u}^{K}=:u$ is uniform convergence due to (3). Thus $u$ is almost periodic (see Corduneanu [1, (1.2)]).
Remark 2.
In some situations, this periodic function $\bar{u}^{K}$ may be different for different choice of $t_{0}$, but we do not go into the detail here.
Acknowledgments. The author is grateful to Professors Chikara Nakayama and Yoshitaka Saiki for valuable comments. Research of TY was partly supported by the JSPS Grants-in-Aid for Scientific Research 20H01819.