### 01 Mar 2024 v4

# Universality of almost periodicity in bounded discrete time series

Abstract.
We consider arbitrary bounded discrete time series. From its statistical feature, without any use of the Fourier transform, we find an almost periodic function which suitably characterizes the corresponding time series.
Key words and phrases:
almost periodic function, chaos in discrete time, recurrent neural network
2020 Mathematics Subject Classification:
Primary 42A20; Secondary 11B50
1. Introduction
In this paper we consider universality of almost periodicity. More precisely, without any use of the Fourier transform, we construct a suitable recursive formula which approximates the corresponding bounded discrete time series in a local time interval, and the corresponding explicit formula tends to an almost periodic function as time tends to infinity.
Let $y:\mathbb{Z}\to[-1,1]$ be a bounded discrete time series arbitrarily chosen, and first let us discretize the range $[-1,1]$ as follows: For $K\in\mathbb{Z}_{\geq 1}$, we choose $\{a_{k}^{K}\}_{k=0}^{K+1}$ such that (for a technical reason, we employ transcendental numbers)
  • $\{a_{1}^{K},a_{2}^{K},\cdots,a^{K}_{K-1},a_{K}^{K}\}\subset\{\pm e^{-\frac{n}{m}}\ ;m,n\in\mathbb{Z}_{\geq 1}\}\subset\mathbb{R}\setminus\mathbb{Q}$,
  • $-1=a_{0}^{K}<a_{1}^{K}<a_{2}^{K}<\cdots<a_{K+1}^{K}=1$,
  • $\sup_{1\leq k\leq K+1}|a_{k-1}^{K}-a_{k}^{K}|\leq^{\exists}\!\!C/K$.
We discretize $y$ as follows:
(1) $\bar{y}_{K}(t):=\operatorname*{arg\,min}_{a\in\{a^{K}_{k}\}_{k=1}^{K}}|y(t)-(a-0)|,$
where $a-0:=a-\varepsilon$ for any sufficiently small $\varepsilon>0$. In what follows we sometimes omit the index $K$ unless confusion occurs. To define a suitable statistical feature, we need to classify patterns in $\bar{y}$. Let $\sigma_{n}$ ($n=1,2,\cdots,N$) be a permutation operator, namely,
$\sigma_{n}:\{1,2,\cdots,L\}\to\{a_{1},a_{2},\cdots,a_{K-1},a_{K}\}$
($\ell\mapsto\sigma_{n}(\ell)$) with $\sigma_{n}\not=\sigma_{n^{\prime}}$ ($n\not=n^{\prime}$), and we impose the following two conditions for determining $N$:
(2) $\begin{cases}\text{For any $t\in\mathbb{Z}$, there is $n\in\{1,\cdots,N\}$ such that}\\ \qquad\text{$\sigma_{n}(\ell)=\bar{y}(t-\ell)$ for $\ell=1,2,\cdots,L$},\\ \text{For any $n\in\{1,\cdots,N\}$ there is $t\in\mathbb{Z}$ such that}\\ \qquad\text{$\sigma_{n}(\ell)=\bar{y}(t-\ell)$ for $\ell=1,2,\cdots,L$.}\end{cases}$
Note that $N\leq K^{L}$ due to the sequence with repetition. The crucial statistical feature is the following: for any $y:\mathbb{Z}\to[-1,1]$, we have the following conditional probability: there exists a $\delta_{K}\in[0,1)$ (of course depending on $K$) such that for any $n\in\{1,2,\cdots,N\}$,
(3) $P\bigg{(}a_{k(n)}=\bar{y}(\cdot)\bigg{|}\sigma_{n}(\ell)=\bar{y}(\cdot-\ell)\quad\text{for}\quad\ell=1,2,\cdots,L\bigg{)}\geq 1-\delta_{K}$
with appropriately chosen $a_{k(n)}$ satisfying the following property:
  • Magnitude relation: For any $K^{\prime}>0$ and $t\in\mathbb{Z}$,
    (4) $\left|a^{K}_{k(n_{t}^{K})}-a^{K+K^{\prime}}_{k(n_{t}^{K+K^{\prime}})}\right|<\frac{C}{K},$
    where $n_{t}^{K}\in\{1,2,\cdots,N\}$ is determined as
    (5) $\bar{y}_{K}(t-\ell)=\sigma_{n_{t}^{K}}(\ell)\quad\text{for}\quad\ell=1,2,\cdots,L.$
Then we can state the main theorem as follows:
Theorem 1.
For any $t_{0}\in\mathbb{Z}$ there exists $u:\mathbb{Z}\to[-1,1]$ such that the following inequality holds:
(6) $\prod_{t^{\prime}=0}^{t}P\bigg{(}|u(t_{0}+t^{\prime})-y(t_{0}+t^{\prime})|<\frac{C}{K}\bigg{)}\geq\left(1-\delta_{K}\right)^{t}$
for any $K\in\mathbb{Z}_{\geq 1}$ and $t\in\mathbb{Z}_{\geq 0}$. Moreover there exists a discretized almost periodic function $ap:\mathbb{Z}\to[-1,1]$ such that
[…]
Remark 1.
Deriving the inequality (6) itself is rather obvious and quite natural due to (3). The nontrivial discovery is explicit construction of $u$ (see (12)), and it tends to the discretized almost periodic function, without any use of Fourier transform.
Corollary 2.
Let us mention the dynamical system case (c.f. [1, 2]). With the aid of Takens’ embedding theorem [6] (see also [5]), it suffices to consider in a delay coordinate. Let
$\begin{split}Y_{t}&:=(y(t),y(t-1),\cdots,y(t-L+1))^{T},\\ \overline{Y}_{t}&:=(\bar{y}(t),\bar{y}(t-1),\cdots,\bar{y}(t-L+1))^{T},\\ \overline{U}_{t}&:=(\bar{u}(t),\bar{u}(t-1),\cdots,\bar{u}(t-L+1))^{T},\end{split}$
where $\bar{u}$ is generated by (12), and let $\Phi:[-1,1]^{L}\to[-1,1]^{L}$ be such that $Y_{t-1}\mapsto Y_{t}$ (in Appendix 3, we give the simplest example of $\Phi$). We see that, for any $t\in\mathbb{Z}_{\geq 0}$, there exists $t^{*}\in\mathbb{Z}_{\geq 0}$ such that
$\begin{cases}\displaystyle\overline{Y}_{t_{0}+t^{*}}=\overline{U}_{t_{0}+t},\quad|Y_{t_{0}+t^{*}}-\overline{U}_{t_{0}+t}|\leq\frac{\sqrt{L}C}{K},\\ Y_{t_{0}+t^{*}}=\Phi(Y_{t_{0}+t^{*}-1}),\\ \displaystyle\overline{Y}_{t_{0}+t^{*}-1}=\overline{U}_{t_{0}+t-1},\quad|Y_{t_{0}+t^{*}-1}-\overline{U}_{t_{0}+t-1}|\leq\frac{\sqrt{L}C}{K}.\end{cases}$
Let
$\beta:=\sup_{t\not=t^{\prime}}|\frac{|\Phi(Y(t))-\Phi(Y(t^{\prime}))|}{|Y(t)-Y(t^{\prime})|}$
and assume $\beta\in\mathbb{R}_{\geq 1}$. By applying the above inequalities inductively, then we have
$\begin{split}&|\bar{u}_{K}(t_{0}+t)-y(t_{0}+t)|\leq|\overline{U}(t_{0}+t)-Y(t_{0}+t)|\\ \leq&|Y(t_{0}+t^{*})-Y(t_{0}+t)|+\frac{\sqrt{L}C}{K}\\ \leq&\beta|Y(t_{0}+t^{*}-1)-Y(t_{0}+t-1)|+\frac{\sqrt{L}C}{K}\\ \leq&\beta|\overline{U}(t_{0}+t-1)-Y(t_{0}+t-1)|+\beta\frac{\sqrt{L}C}{K}+\frac{\sqrt{L}C}{K}\leq\cdots\\ \lesssim&\beta^{t}\left|\overline{U}\left(t_{0}\right)-Y\left(t_{0}\right)\right|+\beta^{t}\frac{\sqrt{L}C}{K}\\ \lesssim&\frac{\sqrt{L}C\beta^{t}}{K}.\end{split}$
This means that
$P\bigg{(}|\bar{u}_{K}(t_{0}+t)-y(t_{0}+t)|\lesssim\frac{\sqrt{L}C\beta^{t}}{K}\bigg{)}=1,$
that is,
[…]
Note that the case $\beta<1$ is trivial, since $\Phi$ becomes a contraction mapping. Again, there exists a discretized almost periodic function $ap:\mathbb{Z}\to[-1,1]$ such that
$|u(t)-ap(t)|\to 0\quad(t\to\infty).$
2. Proof of main theorem
Let $\sigma^{*}_{n}$ ($n=1,2,\cdots,N$) be an adjoint type of permutation operator, more precisely, let
(7) $\sigma^{*}_{n}(\ell-1):=\frac{1}{\sigma_{n}(\ell)}$
for $\ell\in\{1,\cdots,L\}$ and $n\in\{1,\cdots N\}$, where $\sigma^{*}_{i}(0):=\sigma_{i}^{*}(L)$. The definition of the following $N\times N$ matrix $X$ is the key to prove the main theorem:
$X:=\left[h\left(\sum_{\ell=1}^{L}\sigma^{*}_{i}(\ell-1)\sigma_{j}(\ell)\right)\right]_{i,j}$
for some $h:\mathbb{R}\to[-1,1]$,
Lemma 3.
$X$ is a regular matrix provided by appropriately chosen $h:\mathbb{R}\to[-1,1]$.
Proof.
First, by the Lindemann–Weierstrass theorem, we see that
(8) $\frac{a_{k^{\prime}}}{a_{k}}\in(\mathbb{R}\setminus\mathbb{Q})\cup\{-1\}\quad(k\not=k^{\prime}),$
(9) $\sum_{\ell=1}^{L}\frac{a_{k^{\prime}_{\ell}}}{a_{k_{\ell}}}\in(\mathbb{R}\setminus\mathbb{Q})\cup\{-L,-L+1,\cdots,-1,0,1,\cdots,L-1\}$
except for the $k_{1}=k^{\prime}_{1},\ k_{2}=k^{\prime}_{2},\cdots,k_{L}=k^{\prime}_{L}$ case. Also
$\sum_{\ell=1}^{L}\frac{a_{k^{\prime}_{\ell}}^{K}}{a_{k_{\ell}}^{K}}=L$
if and only if $k_{1}=k^{\prime}_{1},\ k_{2}=k^{\prime}_{2},\cdots,k_{L}=k^{\prime}_{L}$. By (7), (8), (9) and Lindemann–Weierstrass theorem, we see
$\sum_{\ell=1}^{L}\sigma^{*}_{j}(\ell-1)\sigma_{j}(\ell)=L\quad\text{and}\quad\sum_{\ell=1}^{L}\sigma^{*}_{i}(\ell-1)\sigma_{j}(\ell)\not=L\quad(i\not=j).$
In order to construct an appropriate $h$, we use a finite set $G$ as follows:
$G:=\left\{\sum_{\ell=1}^{L}\sigma^{*}_{i}(\ell-1)\sigma_{j}(\ell):i,j\in\{1,2,\cdots,N\}\right\}\subset\mathbb{R}.$
Note that $L\in G$. Now we take a smooth function $h:\mathbb{R}\to[-1,1]$ satisfying the following:
$\begin{split}&h(L)\in\mathbb{Q}\setminus\{0\},\\ &h(\gamma)\in\left(\{\pm e^{-\frac{n}{m}}\ ;m,n\in\mathbb{Z}_{\geq 1}\}\cup\{0\}\right)\subset(\mathbb{R}\setminus\mathbb{Q})\cup\{0\},\quad\text{for}\quad\gamma\subset G\setminus\{1\}.\end{split}$
Then we can easily check that (applying the Lindemann–Weierstrass theorem)
  • $h(\gamma_{1})h(\gamma_{2})\cdots h(\gamma_{N})\in(\mathbb{R}\setminus\mathbb{Q})\cup\{0\}$  for  $\{\gamma_{n}\}_{n=1}^{N}\in G^{N}\setminus\{\underbrace{1,1,\cdots,1}_{N}\}$,
  • for any $\{\tau_{n^{\prime}}\}_{n^{\prime}=1}^{N!-1}\in\{-1,1\}^{N!-1}$ and $\{\{\gamma_{n,n^{\prime}}\}_{n=1}^{N}\}_{n^{\prime}=1}^{N!-1}\subset G^{N}\setminus\{\underbrace{1,1,\cdots,1}_{N}\}$,
    $\sum_{n^{\prime}=1}^{N!-1}\tau_{n^{\prime}}h(\gamma_{1,n^{\prime}})h(\gamma_{2,n^{\prime}})\cdots h(\gamma_{N,n^{\prime}})\in(\mathbb{R}\setminus\mathbb{Q})\cup\{0\}.$
By applying the above properties, then we see that the determinant of the matrix $X$ is nonzero, since it is expressed as
$|X|=\eta_{1}+\eta_{2},\quad\eta_{1}\in\mathbb{Q}\setminus\{0\},\quad\eta_{2}\in(\mathbb{R}\setminus\mathbb{Q})\cup\{0\}.$
Now we resume the proof of the main theorem. Let us define $N\times N$ matrices: $Y$ and $W$ as follows:
(10) $Y:=\left[\sum_{\ell=1}^{L}\sigma_{i}^{*}(\ell)\sigma_{j}(\ell)\right]_{i,j}\quad\text{and}\quad W:=YX^{-1}.$
By (2), for each $t\in\mathbb{Z}$, there exists a unique column vector
$e_{n_{t}}:=(\underbrace{\overbrace{0,0,\cdots,0,1}^{n_{t}},0\cdots,0}_{N})^{T}$
such that
(11) $\bar{y}(t-\ell)=\sigma(\ell)e_{n_{t}}\quad\text{for}\quad\ell=1,2,\cdots,L,$
where
$\sigma(\ell):=(\sigma_{1}(\ell),\sigma_{2}(\ell),\cdots,\sigma_{N}(\ell)).$
Thus we have
$W\left(h\left(\sum_{\ell=1}^{L}\sigma^{*}(\ell-1)\bar{y}(t-\ell)\right)\right)=WXe_{n_{t}}=Ye_{n_{t}}=\sum_{\ell=1}^{L}\sigma^{*}(\ell)\bar{y}(t-\ell),$
where $\sigma^{*}(\ell):=(\sigma^{*}_{1}(\ell),\sigma^{*}_{2}(\ell),\cdots,\sigma_{N}^{*}(\ell))^{T}$. To find an appropriate function $u$, we create the explicit formulas $\{\bar{u}(t)\}_{t\geq t_{0}}$ inductively from initial data $\{\bar{y}(t_{0}-\ell)\}_{\ell=1}^{L}$:
(12) $\begin{split}&\begin{cases}\bar{r}(t_{0})&=h(\sum_{\ell=1}^{L}\sigma^{*}(\ell)\bar{y}(t_{0}-\ell)),\\ \bar{u}(t_{0})&=W^{out}\bar{r}(t_{0}),\end{cases}\\ &\begin{cases}\bar{r}(t)&=h(W\bar{r}(t-1)+\sigma^{*}(L)\bar{u}(t-1)-\sigma^{*}(L)\bar{u}(t-L-1)),\\ \bar{u}(t)&=W^{out}\bar{r}(t),\end{cases}\end{split}$
for $t=t_{0}+1,t_{0}+2,\cdots$, where $W^{out}:=AX^{-1}$ and $A:=(a_{k(1)},a_{k(2)},\cdots,a_{k(N)})$. Rigorously we regard $\bar{u}(t-L-1)$ as $\bar{y}(t-L-1)$ for $t=t_{0}+1,t_{0}+2\cdots,t_{0}+L$. Now we show that $\bar{u}_{K}(t)$ tends to a discretized periodic function for any $K$. By (3), for each $K$, we have a chain of the permutation operator $(\sigma_{j},a_{k(j)})$ as follows:
$(\sigma_{j_{1}},a_{k(j_{1})})\rightarrow(\sigma_{j_{2}},a_{k(j_{2})})\rightarrow(\sigma_{j_{3}},a_{k(j_{3})})\rightarrow\cdots.$
Note that the sequence $j_{1},j_{2},j_{3}\cdots$ is different from $n_{1},n_{2},n_{3}\cdots$ in (4). Since the number of permutation operators $\{\sigma_{n}\}_{n=1}^{N}$ is finite, then there exist $m,M\in\mathbb{Z}_{\geq 2}$ ($m<M$) such that $(\sigma_{j_{m}}a_{k(j_{m})})=(\sigma_{j_{M}},a_{k(j_{M})})$. This means that there exists $T_{K}$ such that $\bar{u}_{K}(t)$ is a periodic chain for $t>T_{K}$. Moreover $\lim_{K\to\infty}\bar{u}_{K}=:u$ is uniform convergence due to (4). Thus $u(t)$ tends to an almost periodic function for $t\to\infty$ (see Corduneanu [3, (1.2)]).
Remark 2.
In some situations, this periodic chain may be different for different choice of $t_{0}$, but we do not go into the detail here.
3. Appendix
In this appendix, we give the simplest example of the dynamical system, that is, a linear regression relation. Let $L\in 2\mathbb{Z}_{\geq 1}$. For $\{p_{\ell}\}_{\ell=1}^{L}\subset\mathbb{R}$, we consider the following:
(13) $y(t)=p_{1}y(t-1)+p_{2}y(t-2)+\cdots+p_{L-1}y(t-L+1)+p_{L}y(t-L)\quad(t\geq 0)$
with initial data $\{y(-\ell)\}_{\ell=1}^{L}$. First we figure out the corresponding characteristic equation, that is, we plug the following:
$y(-\ell)=\mu^{L-\ell},\quad(\mu\in\mathbb{R},\ \ell=0,1,\cdots,L)$
into (13). After factorization of the $L$-th degree polynomial, then, we see that $\{p_{\ell}\}_{\ell}$ satisfies the following equality (at least we can easily figure out that $p_{L}=-1$):
$\prod_{j=1}^{L/2}(\mu-e^{i\lambda_{j}})(\mu-e^{-i\lambda_{j}})=\mu^{L}-\sum_{\ell=1}^{L}p_{\ell}\mu^{L-\ell}=0$
if and only if $y$ is a quasi-periodic function:
(14) $y(t)=\sum_{j=1}^{L/2}a_{j}\sin\left(\lambda_{j}(t-b_{j})\right),\quad t\in\mathbb{Z}.$
for $\{a_{j}\}_{j=1}^{L/2}$, $\{b_{j}\}_{j=1}^{L/2}$, $\{\lambda_{j}\}_{j=1}^{L/2}\subset\mathbb{R}$. Otherwise, the time series is not bounded.
Acknowledgments. The author is grateful to Professors Masanobu Inubushi, Chikara Nakayama and Yoshitaka Saiki for valuable comments. Research of TY was partly supported by the JSPS Grants-in-Aid for Scientific Research 20H01819.